Adv. Math. Econ. 5, 55-63 (2003)

Nonlinear generalizations of theorems on inverse–positive matrices

Takao Fujimoto^{1}, José A. Silva^{2} and Antonio Villar^{3}

1 Department of Economic Analysis, University of Alicante, 03071 Alicante, Spain

Department of Economics, Kagawa University, Takamatsu, Kagawa, 760, Japan 2 Department of Economic Analysis, University of Alicante, 03071 Alicante, Spain 3 Department of Economic Analysis, University of Alicante and IVIE, 03071 Ali-

cante, Spain

Received: August 2, 2002 Revised: September 2, 2002 JEL classification: C67 Mathematical Subject Classification (2000): 15A48 Abstract. Nonlinear extensions of some theorems concerning inverse-positive matrices are made. The method of proof is based on an intuitively natural geometrical consideration.